Circular No. 6199

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET P/1995 O2 (MUELLER 1) J. V. Scotti, Lunar and Planetary Laboratory, reports his recovery of

comet P/1987 U2 (= 1987 XXXI = 1987a_{1}) with the 0.9-m Spacewatch telescope at Kitt Peak. The comet’s image is very faint and very slightly diffuse, near the limit. The indicated correction (ignoring perturbations) to the prediction by S. Nakano on MPC 22030 is ΔT = −0.75 day.

1995 UT  α_{2000}  δ_{2000}  m_{1}
July 30.43941  23^{h}14^{m}25.^{s}45  −11^{o}43′36′′.0  22.2
30.45900  23 14 25.07  −11 43 36.8  21.9
30.46644  23 14 25.01  −11 43 37.8  22.5
Aug. 1.42770  23 13 50.27  −11 45 55.9  21.9
1.44590  23 13 49.91  −11 45 58.0
1.46213  23 13 49.57  −11 45 57.6

The following improved orbital elements satisfy 34 observations 1987– 1995 with mean residual 0′′.70:

Epoch = 1996 Apr. 27.0 TT

T = 1996 Apr. 24.6661 TT  ω = 29.^{o}9207
e = 0.337378  Ω = 4.5613 }2000.0
q = 2.739482 AU  i = 8.7956

a = 4.134308 AU n^{o} = 0.1172465 P = 8.406 years

COMET 32P/COMAS SOLÁ Scotti also provides Spacewatch observations of this comet. On Aug.

1.44 UT, m_{2} = 21.9, there was a coma diameter of 6′′ and a tail 0′.60 long in p.a. 254^{o}. On Aug. 2.4, m_{1} = 19.7–20.0.

SUPERNOVA 1995U IN ESO 235-IG13 S. Benetti, European Southern Observatory, writes: “I have observed

SN 1995U on Aug. 3.12 UT with the Danish 1.54-m telescope (+ DFOSC) at La Silla. The spectrum (range 396–793 nm, resolution 1 nm) classifies SN 1995U as a type Ia supernova, ∼ 3 weeks past maximum light. Broad emissions are measured at 418.0, 451.0, 468.0, 521.6, 558.3, 619.5 and 677.2 nm. The expansion velocity deduced from the minimum of the Si II 635.5nm line is ∼ 9200 km/s (the redshift of parent galaxy, determined from Mg I and Na I absorption lines, being 0.0556).”

1995 August 3 Brian G. Marsden